# DiffBind: An SE(3) Equivariant Network for Accurate All-Atom Protein-Ligand Docking with Full Pocket Side Chain Flexibility | Jintao Zhu, Zhonghui Gu, Jianfeng Pei, and Luhua Lai |

### 15 Dec 2023 v2

@@ 2.2 Performance on the Posebusters test set
To evaluate the physical plausibility of poses generated by DiffBind, we compared its performance to other baseline methods using the Posebusters test set and the Posebusters suite 47, a tool designed to assess the validity of ligand-protein complexes based on criteria including bond length, planarity of aromatic rings in ligands, and clashes between ligands and proteins. Success in docking is redefined as a pose having an L-RMSD less than 2 Å and simultaneously passing the physical validity check by Posebusters, with this success rate termed as the PB-success rate. The Posebusters test set comprises 428 distinct complexes released since 2021, with no overlap with the PDBBind v2020 dataset. To demonstrate that the success rate of DiffBind is not solely due to local ligand energy relaxation and its superior in side chain packing, the performance of other methods is evaluated with stricter energy minimization for the ligand, given the experimental side chain conformation. For poses generated by DiffBind, ligand energy minimization is conducted using the side chains as predicted by the model. The energy minimization is performed using the AMBER ff14sb force field 59 for prorein and the Sage force field 60 for ligand in OpenMM 61, as used in the Posebusters paper 47. Fig.4.(A) shows that traditional rigid receptor docking methods like Glide perform best on re-docking when provided with the correct Holo pocket environment, followed by Vina and Gold, with most of their generated docking poses being physically valid. However, their performance significantly deteriorates when docking with Rosetta-repacked proteins, highlighting their heavy dependence on side chain conformations. Traditional flexible docking methods rDock and VinaFlex are also involved in comparison. VinaFlex, heavily reliant on predefined flexible side chains, performs the worst in our scenario where information about flexible side chains is assumed unavailable. rDock, capable of optimizing functional groups prone to forming hydrogen bonds in side chains, achieves higher success rate in repacked proteins compared to Vina and Glide, but lower success rate in proteins with ground-truth side chains. For these traditional methods, their PB-success rate is only slightly lower than their overall success rate, indicating that most generated poses is validated by Posebusters suite due to the physical components in their scoring functions. Therefore, the post ligand optimization using force field does not cause obvious impact to their PB-success rate. Among blind docking methods, DiffDock shows better performance (success rate of 38) than TankBind (16) and EquiBind (2), but most of their generated poses are invalid due to ignoring protein side chains (Fig.4.(A)). Ligand energy minimization significantly improves the PB-success rate of DiffDock (35). TankBind and EquiBind also see improvements in PB-success rate with energy minimization, but still lag behind DiffDock (Fig.4.(B)). Although blind docking is a tough task for its broad searching space in the whole protein, but flexible docking like DiffBind, denoising a chaotic side chain conformation into a well-packed conformation having valid interaction with the ligand, has much more objectives for prediction. DiffBind, utilizing Smina scoring function or MDN network as the confidence model to select the top-1 pose from 40 generated ones, outperforms all other deep learning-based methods in both blind and pocket docking. DiffBind-Smina and DiffBind-MDN demonstrate both high success rate (50.2 for DiffBind-Smina and 48.1 for DiffBind-MDN) and PB-success rate (49.1 for DiffBind-Smina and 44.4 for DiffBind-MDN), with lower penalties by Posebusters compared to other deep learning-based methods, showcasing the capability of DiffBind in generating accurate and physically plausible complex poses. The performance of DiffBind is comparable to traditional rigid receptor docking methods using known ground-truth side chain conformations for redocking. As is depicted from Supplemental Fig. S6, DiffBind shows its effectiveness in binding site identification and pocket side chain recovery on Posebusters test set, as well. Force field optimization has minimal impact on DiffBind generated structures, which also demonstrates the high physical plausibility of DiffBind generated poses. Among other deep learning-based pocket docking methods, KarmaDock Align achieves the highest success rate (30.4) but a very low PB-success rate (6.1). Force field optimization of ligands rescues most poses with L-RMSD 2 Å into good physical validity. KarmaDock Align, EDM-Dock, TankBind-pocket, Uni-Mol, and EDM-Dock, which focus on fitting the RMSD of the ligand during training, tend to ignore the intra energy of the generated poses and protein side chains, as is shown from supplemental Fig. S8. Indeed, force field optimization is not allowed in realistic docking to meet the demands of high-throughput screening.
@@
Four specific cases from the Posebusters test set (supplementary Fig. S9), never trained or seen by DiffBind, are presented to highlight its superiority over other methods focusing solely on ligand coordinates while neglecting ligand conformation validity. In complexes with PDB ID 6TW5, 7PRM, 7T1D, and 7CD9, DiffBind successfully docks ligands into precise positions with valid conformations and recovers pocket side chains into good interaction with ligands. In contrast, KarmaDock Align, EDM-Dock, and TankBind-pocket fails to predict correct binding ligand poses, and their generated poses cannot pass the physical plausibility check of the Posebusters suite. As is shown from Supplementary Table. S4, ligand poses generated by EDM-Dock and TankBind-pocket exhibit both internal invalidity (including internal steric clash, bump aromatic ring, etc) and steric clash with proteins, while KarmaDock Align, due to using RDKit for ligand pose alignment, frequently fails in ligand-protein clash.
| Method | Ensemble-CDK2 | Ensemble-EGFR | Ensemble-FXA | ApoRef | CASF2016 | GPCR-AF2 | |||||||||||||
RMSD Mean / Å |
RMSD Median / Å |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
RMSD Mean |
RMSD Median |
PB-SR |
Traditional rigid receptor docking methods |
Vina |
6.12±2.63 |
6.26 |
0.079 |
6.37±2.34 |
6.90 |
0.060 |
4.35±3.17 |
3.02 |
0.344 |
6.98±3.35 |
7.22 |
0.082 |
4.93±3.62 |
4.46 |
0.294 |
7.42±3.99 |
7.77 |
0.136 |
| LinF9 |
6.12±2.50 |
6.31 |
0.061 |
6.31±2.43 |
6.63 |
0.119 |
3.88±2.90 |
2.96 |
0.365 |
6.59±3.08 |
6.67 |
0.089 |
4.99±3.56 |
4.52 |
0.272 |
5.66±2.54 |
5.18 |
0.076 |
| Smina |
6.15±2.65 |
6.28 |
0.079 |
6.33±2.53 |
6.90 |
0.090 |
4.17±3.16 |
2.70 |
0.362 |
7.08±3.47 |
7.43 |
0.077 |
4.88±3.69 |
4.38 |
0.304 |
7.25±4.15 |
7.41 |
0.152 |
| Gnina |
5.67±2.63 |
5.78 |
0.099 |
6.02±2.69 |
6.29 |
0.090 |
3.83±2.99 |
2.44 |
0.394 |
6.67±3.65 |
6.51 |
0.082 |
4.57±3.56 |
3.74 |
0.320 |
6.72±3.95 |
6.06 |
0.106 |
| Glide |
4.81±2.60 |
5.21 |
0.154 |
6.87±3.63 |
7.67 |
0.090 |
3.49±2.72 |
2.31 |
0.271 |
6.62±3.51 |
6.85 |
0.091 |
4.28±3.20 |
3.74 |
0.219 |
5.37±3.63 |
4.57 |
0.182 |
Traditional flexible docking methods |
VinaFlex |
8.04±2.92 |
7.88 |
0.013 |
9.03±2.47 |
9.446 |
0.000 |
9.19±2.22 |
9.42 |
0.005 |
8.81±2.94 |
9.06 |
0.015 |
8.80±2.98 |
9.07 |
0.023 |
7.36±3.33 |
7.07 |
0.045 |
| rDock |
4.61±2.76 |
4.62 |
0.257 |
5.45±2.98 |
4.70 |
0.134 |
3.29±2.66 |
2.28 |
0.440 |
5.46±3.02 |
5.31 |
0.157 |
4.11±2.82 |
3.58 |
0.296 |
5.37±3.60 |
5.12 |
0.212 |
Deep learning-based docking methods |
TankBind-pocket |
2.17±1.85 |
1.62 |
0.100 |
2.21±0.96 |
1.82 |
0.015 |
1.73±0.96 |
1.50 |
0.110 |
2.63±1.88 |
1.97 |
0.040 |
2.37±2.38 |
1.63 |
0.123 |
3.74±1.98 |
3.42 |
0.015 |
| EDM-Dock |
2.62±1.27 |
2.32 |
0.051 |
2.95±1.05 |
2.66 |
0.015 |
2.68±0.98 |
0.45 |
0.009 |
3.33±1.55 |
3.09 |
0.011 |
3.06±1.92 |
2.55 |
0.064 |
4.72±1.95 |
4.61 |
0.00 |
| KarmaDock Align |
1.89±1.20 |
1.58 |
0.135 |
2.65±2.06 |
1.93 |
0.045 |
2.43±0.92 |
2.30 |
0.009 |
2.47±1.56 |
2.04 |
0.047 |
2.36±1.57 |
1.89 |
0.136 |
4.30±2.20 |
3.91 |
0.015 |
| DiffBind-Smina |
2.31±1.80 |
1.73 |
0.564 |
3.44±2.53 |
2.57 |
0.403 |
1.54±1.14 |
1.30 |
0.789 |
2.96±2.32 |
2.10 |
0.434 |
2.30±2.28 |
1.58 |
0.566 |
4.92±4.12 |
4.02 |
0.227 | |
| DiffBind-MDN |
1.85±1.34 |
1.48 | 0.674 |
2.58±2.15 |
1.80 | 0.478 |
1.42±0.70 |
1.35 |
0.794 |
2.32±1.73 |
1.76 | 0.476 |
1.87±1.93 |
1.25 | 0.636 |
3.64±2.59 |
2.74 |
0.212 |
Best performance in bold. RMSD Mean and RMSD Medium, lowest; PB-SR, highest. RMSD Mean and RMSD Median denote the average ± standard deviation and median of L-RMSD for top-1 generated ligand poses from each complex, respectively. PB-SR denotes PB-success rate.

### 19 Dec 2023 v3

@@ 4.2 Performance on the Posebusters test set
Given that the similarity between samples to be predicted and those used in training can influence the performance of deep learning methods, Buttenschoen et al. 47, aiming for a more equitable comparison with traditional docking methods based on scoring functions, have curated a dataset called Posebusters test set from the PDB database. The Posebusters test set exclusively comprises 428 complexes on which the deep learning methods have not been trained. To evaluate the physical plausibility of poses generated by DiffBindFR, we compared its performance to other baseline methods using the Posebusters test set and the Posebusters suite 47, a tool designed to assess the validity of ligand-protein complexes based on criteria including bond length, planarity of aromatic rings in ligands, and clashes between ligands and proteins. Success in docking is redefined as a pose having an L-RMSD less than 2 Å and simultaneously passing the physical validity check by Posebusters, with this success rate termed as the PB-success rate. The Posebusters test set comprises 428 distinct complexes released since 2021, with no overlap with the PDBbind v2020 dataset. To demonstrate that the success rate of DiffBindFR is not solely due to local ligand energy relaxation and its superior in side chain packing, the performance of other methods is evaluated with stricter energy minimization for the ligand, given the experimental side chain conformation. For poses generated by DiffBindFR, ligand energy minimization is conducted using the side chains as predicted by the model. The energy minimization is performed using the AMBER ff14sb force field 69 for prorein and the Sage force field 70 for ligand in OpenMM 71, as used in the Posebusters paper 47. Fig.4.(A) shows that traditional rigid receptor docking methods like Glide perform best on re-docking when provided with the correct Holo pocket environment, followed by Vina and Gold, with most of their generated docking poses being physically valid. However, their performance significantly deteriorates when docking with Rosetta-repacked proteins, highlighting their heavy dependence on side chain conformations. Traditional flexible docking methods rDock and VinaFlex are also involved in comparison. VinaFlex, heavily reliant on predefined flexible side chains, performs the worst in our scenario where information about flexible side chains is assumed unavailable. rDock, capable of optimizing functional groups prone to forming hydrogen bonds in side chains, achieves higher success rate in repacked proteins compared to Vina and Glide, but lower success rate in proteins with ground-truth side chains. For these traditional methods, their PB-success rate is only slightly lower than their overall success rate, indicating that most generated poses is validated by Posebusters suite due to the physical components in their scoring functions. Therefore, the post ligand optimization using force field does not cause obvious impact to their PB-success rate.
| Method |
PB-success rate |
| Ensemble-CDK2 | Ensemble-EGFR | Ensemble-FXA | ApoRef | CASF2016 | GPCR-AF2 | ||
| Traditional rigid receptor docking methods | Vina | 0.079 | 0.060 | 0.344 | 0.082 | 0.294 | 0.136 |
| LinF9 | 0.061 | 0.119 | 0.365 | 0.089 | 0.272 | 0.076 | |
| Smina | 0.079 | 0.090 | 0.362 | 0.077 | 0.304 | 0.152 | |
| Gnina | 0.099 | 0.090 | 0.394 | 0.082 | 0.320 | 0.106 | |
| Glide | 0.154 | 0.090 | 0.271 | 0.091 | 0.219 | 0.182 | |
| Traditional flexible docking methods | VinaFlex | 0.013 | 0.000 | 0.005 | 0.015 | 0.023 | 0.045 |
| rDock | 0.257 | 0.134 | 0.440 | 0.157 | 0.296 | 0.212 | |
| Deep learning-based docking methods | TankBind-pocket | 0.100 | 0.015 | 0.110 | 0.040 | 0.123 | 0.015 |
| EDM-Dock | 0.051 | 0.015 | 0.009 | 0.011 | 0.064 | 0.00 | |
| KarmaDock Align | 0.135 | 0.045 | 0.009 | 0.047 | 0.136 | 0.015 | |
| DiffBindFR-Smina | 0.564 | 0.403 | 0.789 | 0.434 | 0.566 | 0.227 | |
| DiffBindFR-MDN | 0.674 | 0.478 | 0.794 | 0.476 | 0.636 | 0.212 | |
Best performance in bold for the highest PB-success rate.
Among blind docking methods, DiffDock shows better performance (success rate of 38) than TankBind (16) and EquiBind (2), but most of their generated poses are invalid due to ignoring protein side chains (Fig.4.(A)). Ligand energy minimization significantly improves the PB-success rate of DiffDock (35). TankBind and EquiBind also see improvements in PB-success rate with energy minimization, but still lag behind DiffDock (Fig.4.(B)). Although blind docking is a tough task for its broad searching space in the whole protein, flexible pocket docking method like DiffBindFR, denoising a chaotic side chain conformation into a well-packed conformation having valid interaction with the ligand, has much more objectives for prediction. DiffBindFR, utilizing Smina scoring function or MDN network as the confidence model to select the top-1 pose from 40 generated ones, outperforms all other deep learning-based blind docking methods and pocket docking methods. DiffBindFR-Smina and DiffBindFR-MDN demonstrate both high success rate (50.2 for DiffBindFR-Smina and 48.1 for DiffBindFR-MDN) and PB-success rate (49.1 for DiffBindFR-Smina and 44.4 for DiffBindFR-MDN), with lower penalties by Posebusters compared to other deep learning-based methods, showcasing the capability of DiffBindFR in generating accurate and physically plausible complex poses. The performance of DiffBindFR is comparable to traditional rigid receptor docking methods using known ground-truth side chain conformations for redocking. As is depicted from Supplemental Fig. S7, DiffBindFR shows its effectiveness in binding site identification and pocket side chain recovery on Posebusters test set, as well. Force field optimization has minimal impact on DiffBindFR generated structures, which also demonstrates the high physical plausibility of DiffBindFR generated poses. Among other deep learning-based pocket docking methods, KarmaDock Align achieves the highest success rate (30.4) but a very low PB-success rate (6.1). Force field optimization of ligands rescues most poses with L-RMSD 2 Å into good physical validity. KarmaDock Align, EDM-Dock, TankBind-pocket, Uni-Mol, and EDM-Dock, which focus on fitting the RMSD of the ligand during training, tend to ignore the intra energy of the generated poses and protein side chains, as is shown from supplemental Fig. S9. Indeed, force field optimization is not allowed in realistic docking to meet the demands of high-throughput screening.
@@
Four specific cases from the Posebusters test set (supplementary Fig. S10), never trained or seen by DiffBindFR, are presented to highlight its superiority over other methods focusing solely on ligand coordinates while neglecting ligand conformation validity. In complexes with PDB ID 6TW5, 7PRM, 7T1D, and 7CD9, DiffBindFR successfully docks ligands into precise positions with valid conformations and recovers pocket side chains into good interaction with ligands. In contrast, KarmaDock Align, EDM-Dock, and TankBind-pocket fail to predict correct binding ligand poses, and their generated poses cannot pass the physical plausibility check of the Posebusters suite. As is shown in Supplementary Table. S4, ligand poses generated by EDM-Dock and TankBind-pocket exhibit both internal invalidity (including internal steric clash, bump aromatic ring, etc) and steric clash with proteins, while KarmaDock Align, due to using RDKit for ligand pose alignment, frequently fails in reducing ligand-protein clash.